# Universal control of four singlet-triplet qubits

### v2

@@ III Universal single-qubit control
Figure 3: Two-qubit interactions across the quantum dot ladder. a, A plaquette of two connected DQDs. The $S-T_{-}$ qubits have a splitting of $J_{ij}-\overline{E}_{z,ij}$ and $J_{kl}-\overline{E}_{z,kl}$ (neglecting $\Delta_{ST_{-}}$), which are controlled by the detunings $\varepsilon_{ij}$ and $\varepsilon_{kl}$, respectivily. The qubit-qubit coupling $J_{\text{coup}}$ is an average of $J_{ik}$ and $J_{jl}$ between the corresponding dots, which are controlled by $\varepsilon_{ik}$ and $\varepsilon_{jl}$. b, The energy levels of two-qubit states, where we fix $\varepsilon_{kl}$ to be positively large and scan $\varepsilon_{ij}$. At the positions where $J_{ij}-\overline{E}_{z,ij}$ equals $J_{kl}-\overline{E}_{z,kl}$, an anticrossing with a gap $J_{\text{coup}}$ forms (black dashed circles), which can be used to induce SWAP oscillations between |$\ket{ST_{-}}$ and |$\ket{T_{-}S}$. The parameters we use in this calculation are based on the experimental results for Q3-Q4 shown in panel f. c, The pulse scheme for SWAP operations. We start in (0,2) or (2,0), at large positive or negative detuning, and diabatically pulse one qubit to (1,1) at modest detuning such that it remains in |$\ket{S}$, and pulse the other qubit to zero detuning where a $\pi$ rotation for a time $t_{\pi}$ takes it to |$\ket{T_{-}}$. At this point, the qubits are set to |$\ket{ST_{-}}$ or |$\ket{T_{-}S}$. Next, we pulse the detunings of both qubits to make their energies resonant, while at the same time activating $J_{ik}$ and $J_{jl}$. This will kickstart SWAP oscillations between the two qubits. The dashed lines in the pulse of $\varepsilon_{ij}$ show that we scan the detuning of one qubit to find the condition for SWAP operations. After an evolution time $t_{\text{wait}}$, we pulse the detunings to the PSB readout configuration for one of the qubits. d-f, The experimental results of SWAP oscillations, showing measured triplet probabilities $P_{T}$ or singlet probabilities $P_{S}$ as a function of operation time $t_{\text{wait}}$ and the detuning voltage for Q1-Q2 (d), Q2-Q3 (e) and Q3-Q4 (f). The initial states of two qubits (before the SWAP oscillations) are denoted on the top, and the qubit pair that is read out is indicated by the dashed arrow showing the readout pulse direction. g, Measured singlet probabilities $P_{S}$ as a function of the evolution time $t_{\text{wait}}$ at the center of the chevron patterns of the SWAP oscillations for each pair of qubits, with a detuning voltage $\varepsilon_{15}=-6$ mV, $\varepsilon_{26}=2.2$ mV, and $\varepsilon_{48}=6.2$ mV. The conditions are slightly different from the data in d-f due to device tuning.
The measured $z$-axis rotations as a function of the change of barrier voltage $\delta\text{v}B_{ij}$ for the four qubits are shown in Fig. 2g-j. The oscillation frequency is given by $f_{ST_{-}}=\sqrt{(J-\overline{E}_{z})^{2}+\Delta^{2}_{ST_{-}}}/h$, where $h$ is Planck’s constant. Fig. 2n summarizes how $f_{ST_{-}}$ can be tuned via $\delta\text{v}B_{ij}$. We note that the outer two barrier gates $\text{v}B_{15}$ and $\text{v}B_{48}$ have a stronger effect on the corresponding $J_{ij}$ than the inner barrier gates $\text{v}B_{26}$ and $\text{v}B_{37}$. This may be explained by additional residues below the inner barrier gates, which are fabricated in the last step [38], and by the different fan-out routing for the outer barrier gates (see Fig. 1a,b). With the tuning range shown in panel n, the highest ratio $(J-\overline{E}_{z})/\Delta_{ST_{-}}$ amounts to around 20 for the outer qubits Q1 and Q4 and about 10 for the inner qubits Q2 and Q3. When the external magnetic field strength is varied while keeping the gate voltages fixed, the oscillation speed increases linearly with the field due to the contribution from Zeeman energy in $f_{ST_{-}}$ (see Fig. 2m). From the slope, we extract $\overline{g}_{ij}$ for the four qubits, obtaining values of 0.33(1) - 0.37(1), similar to previous devices [10].
The dephasing time under free evolution, which is traditionally termed $T_{2}^{*}$, is an important metric for assessing the qubit quality. Since the qubit rotations around both the $x$-axis and the $z$-axis are the result of free evolution, we here introduce $T_{x}^{*}$ and $T_{z}^{*}$ to describe the corresponding dephasing times. Fig. 2o,p show the measured damped oscillations of the qubits under $x$-axis and $z$-axis control, to which we fit with a function $P_{T}=P_{0}+A\text{cos}(2\pi ft+\phi)\text{exp}[-(t/T^{*})^{\beta}]$. $P_{0}$, $A$, $\beta$, $f$, $T^{*}$ are fitting parameters, where $f$ refers to the oscillation frequency and $T^{*}$ refers to $T_{x}^{*}$ or $T_{z}^{*}$. From the fits, we obtain a $T_{x}^{*}$ of 0.5 - 2.1 $\mu$s and a $T_{z}^{*}$ of 42(5) - 147(31) ns. The measured values of $T_{x}^{*}$ are slightly lower than previously reported values measured at $B=3$ mT [19] under the same condition that $B$ is parallel to the hole movement direction. This can be partly attributed to the larger magnetic field in the present experiment: $\Delta_{ST_{-}}$ scales with $B$ [44, 45] and not only sets $f_{ST_{-}}$ but also constitutes a proportionality factor between noise and $f_{ST_{-}}$ fluctuations (see also Supplementary Information section VIII). The extracted $x$-axis rotation frequencies in Fig. 2o reflect the values of $\Delta_{ST_{-}}$ for each qubit, which are around 11.9-15.9 MHz, nearly an order of magnitude larger than the results reported at $B=3$ mT [19]. This confirms that $\Delta_{ST_{-}}$ is stronger in the present experiment than in the previous work at 3 mT. The small variation in $\Delta_{ST_{-}}$ and in the measured average $g$-factors suggests a homogeneous spin-orbit coupling in this device. In comparison, the variations in $T_{x}^{*}$ are quite large, which may be caused by spatially dependent charge noise or inhomogeneous hyperfine-induced dephasing due to dot size differences [55].
We also observe that $T_{z}^{*}$ is nearly an order of magnitude smaller than $T_{x}^{*}$. Possibly this is due to the fact that fluctuations in the tunnel barrier translate to fluctuations in $J$, which couple in directly during $z$-axis evolution but only to second order for $x$-axis evolution [56, 57]. Additionally, the increased curvature of the singlet branch for larger $J$ implies a higher sensitivity to detuning noise. The variations in $J$ thus may contribute to the spread of the $T_{z}^{*}$ values we obtained in Fig. 2p.
@@ IV Two-qubit gate
where $ij$ and $kl$ refer to the respective qubit dot pair, and the interqubit coupling $J_{\text{coup}}=(J_{ik}+J_{jl})/2$. The coupling term is reminiscent of two well-known interaction Hamiltonians. If the factor $1/2$ of the $\sigma_{z}\sigma_{z}$ coupling term were 1 instead, we recover the exchange Hamiltonian that generates the SWAP gate and the universal $\sqrt{\text{SWAP}}$ gate. If that factor were zero, only the flip-flop terms would survive, which generate the iSWAP and $\sqrt{\text{iSWAP}}$ gate. The coupling Hamiltonian in Eq. 2 thus generates a SWAP-style gate that is not a standard two-qubit gate but is also universal from the perspective of quantum computing (see Supplementary Information section VIII). For simplicity, we call it a SWAP gate in the remainder of this work.
@@
Fig. 3d-f show the resulting SWAP oscillations for Q1-Q2, Q2-Q3, and Q3-Q4. Chevron-style patterns are observed with the energies of the two qubits aligned in the middle of the patterns. Moving away from the middle, the energy of one qubit is shifted with respect to that of the other. This qubit-qubit energy detuning tilts the rotation axis and accelerates the rotation. Looking closely, the chevron patterns are not symmetrical. This can be understood by the fact that the qubit energy does not vary linearly with detuning. In some panels, single-qubit oscillations around an axis close to the $x$-axis are also observed, such as the data at $\varepsilon_{15}=-2$ mV in Fig. 3d and that at $\varepsilon_{48}=2.5$ mV in Fig. 3f. These $\varepsilon_{ij}$ values are close to zero interdot detuning, and when $J(\varepsilon_{ij}=0)$ is not much larger than $\overline{E}_{z}$, such rotations are expected. We note that SWAP oscillations between |$\ket{ST_{-}}$ and |$\ket{T_{-}S}$ were also observed in previous research on simulating the dynamics of an antiferromagnetic spin chain and resonating-valence-bond states based on the Heisenberg model in four-quantum-dot systems [18, 19].
The oscillation frequencies in the middle of the chevron patterns are in the range of 22.2(3) - 112(1) MHz, corresponding to a $\sqrt{\text{SWAP}}$ (entangling) gate with durations of just 2.2 ns to 11.3 ns, more than an order of magnitude faster than the entangling gate based on capacitive coupling [13, 14]. To determine the dephasing times of the SWAP oscillations, we collect data in the middle of the chevron patterns, as shown in Fig. 3g, and fit them with the same function as used for single-qubit oscillations. The extracted dephasing times are between 63(24) and 154(43) ns. The dephasing times may be further increased by executing the SWAP oscillations at the symmetry points of the detuning of each qubit, which requires a stronger tunability of the exchange interactions using the barrier gates.
@@
Figure 4: Implementation of a quantum circuit for entanglement generation and quantum state transfer. a, Quantum circuit with all the qubits initialized into the singlet. A $\pi$ rotation around the $x$-axis (32 ns) rotates Q2 into a triplet state, and a SWAP interaction for a variable time $t_{\text{SWAP}}$ periodically produces entanglement between Q1 and Q2. Two subsequent SWAP gates (30 ns and 24 ns) transfer the state of Q2 to Q4 and a final single-qubit rotation of Q4 for a variable time $t_{\text{Q4}}$ is followed by the Q4 readout. The delay time between each quantum gate is set to zero. b,c, Experimental (b) and numerical (c) results after running the quantum circuit of panel a, with triplet probabilities $P_{T}$ of Q4 shown as a function of $t_{\text{SWAP}}$ and $t_{\text{Q4}}$. d, Linecuts from a showing triplet probabilities $P_{T}$ of Q4 as a function of control time $t_{\text{Q4}}$. The data is vertically shifted for clarity.
@@ V Quantum circuit implementation
Finally, using a combination of the single- and two-qubit gates demonstrated above, we aim to implement a quantum circuit designed to create and distribute an entangled state. As shown in Fig. 4a, we first initialize Q1 and Q2 into |$\ket{ST_{-}}$ by applying a $\pi$ rotation on Q2 and then activate a SWAP interaction between Q1 and Q2 for a duration $t_{\text{SWAP}}$. This interaction is expected to generate entanglement when $t_{\text{SWAP}}$ corresponds to a quarter period, i.e. for a $\sqrt{\text{SWAP}}$ gate. Next, we apply consecutive half-period SWAP gates of Q2-Q3 and Q3-Q4 to transfer the state of Q2 to Q4 via Q3. Finally, we perform a single-qubit $x$-axis rotation of Q4 for a time $t_{\text{Q4}}$ and measure Q4.
The experimental results are shown in Fig. 4b, where the single-qubit oscillations of Q4 as a function of $t_{\text{Q4}}$ are modulated in phase by $t_{\text{SWAP}}$, resulting in a checkerboard pattern. The underlying mechanism is that the state of Q2 oscillates as a function of $t_{\text{SWAP}}$, as quantum information is periodically exchanged between Q1 and Q2. Therefore the state of Q4 following the quantum state transfer also oscillates with $t_{\text{SWAP}}$. Where the evolution of Q4 changes phase, $t_{\text{SWAP}}$ corresponds to the duration of a $\sqrt{\text{SWAP}}$ operation (modulo an integer number of SWAP operations), at which point maximal entanglement between Q1 and Q2 is expected. When two qubits are maximally entangled, the density matrix of each qubit by itself is fully mixed. At this point, the measured $P_{T}$ of Q4 should not oscillate as a function of $t_{\text{Q4}}$. This is indeed what we observe, see Fig. 4d, where we show the linecuts from Fig. 4b. A trace without oscillations is observed between two sets of out-of-phase oscillations of Q4, as expected. The same features are seen in Fig. 4c, which shows the ideally expected checkerboard pattern obtained from numerical simulations of the protocol of Fig. 4a, assuming perfect initialization, operations and readout.
@@ VI Conclusion
In conclusion, we have experimentally demonstrated initialization, readout, and universal control of four singlet-triplet ($S-T_{-}$) qubits in a 2x4 germanium quantum dot array. The intrinsic spin-orbit coupling in germanium combined with tunnel barrier control of the exchange interactions within each qubit, allows us to realize two-axis control of the four individual $S-T_{-}$ qubits using baseband signals only. Furthermore, through independent control of the exchange interactions between any pair of neighbouring spins across the device, we are able to demonstrate $\sqrt{\text{SWAP}}$ and SWAP gates for each neighboring pair of qubits and implement a quantum circuit that spans the entire array. With four universally controlled qubits in a bilinear array, these results put baseband-controlled singlet-triplet spin qubits in germanium firmly on the map as a potential candidate for large-scale quantum computing.
In future experiments, the control fidelities and error channels of these qubits can be thoroughly characterized by various metrology methods such as randomized benchmarking and gate set tomography [7, 15]. The qubit quality and control fidelity can be potentially improved by tuning the qubits into the operational sweet spot with respect to the electric field or magnetic field [58, 59, 55], optimizing control pulses [28, 47], reducing charge noise [60], further improvements in the Ge/SiGe heterostructure [61, 62], and suppressing nuclear spin noise using substrates with reduced spinful nuclei [41]. Moreover, with programmable control of exchange interactions in the array, this spin ladder can also be used for analog simulation of a wealth of rich physical phenomena such as quantum magnetism [63].

### v3

@@ III Universal single-qubit control
Figure 3: Two-qubit interactions across the quantum dot ladder. a, A plaquette of two connected DQDs. The $S-T_{-}$ qubits have a splitting of $J_{ij}-\overline{E}_{z,ij}$ and $J_{kl}-\overline{E}_{z,kl}$ (neglecting $\Delta_{ST_{-}}$), which are controlled by the detunings $\varepsilon_{ij}$ and $\varepsilon_{kl}$, respectivily. The qubit-qubit coupling $J_{\text{coup}}$ is an average of $J_{ik}$ and $J_{jl}$ between the corresponding dots, which are controlled by $\varepsilon_{ik}$ and $\varepsilon_{jl}$. b, The energy levels of two-qubit states, where we fix $\varepsilon_{kl}$ to be positively large and scan $\varepsilon_{ij}$. At the positions where $J_{ij}-\overline{E}_{z,ij}$ equals $J_{kl}-\overline{E}_{z,kl}$, an anticrossing with a gap $J_{\text{coup}}$ forms (black dashed circles), which can be used to induce SWAP oscillations between |$\ket{ST_{-}}$ and |$\ket{T_{-}S}$. The parameters used in this calculation are based on the experimental results for Q3-Q4 shown in Supplementary Fig. 4c. c, The pulse scheme for SWAP operations. We start in (0,2) or (2,0), at large positive or negative detuning, and diabatically pulse one qubit to (1,1) at modest detuning such that it remains in |$\ket{S}$, and pulse the other qubit to zero detuning where a $\pi$ rotation for a time $t_{X}$ takes it to |$\ket{T_{-}}$. At this point, the qubits are set to |$\ket{ST_{-}}$ or |$\ket{T_{-}S}$. Next, we pulse the detunings of both qubits to make their energies resonant, while at the same time activating $J_{ik}$ and $J_{jl}$. This will kickstart SWAP oscillations between the two qubits. The dashed lines in the pulse of $\varepsilon_{ij}$ show that we scan the detuning of one qubit to find the condition for SWAP operations. After an evolution time $t_{\text{wait}}$, we pulse the detunings to the PSB readout configuration for one of the qubits. d-f, The experimental results of SWAP oscillations, showing measured triplet probabilities $P_{T}$ or singlet probabilities $P_{S}$ as a function of operation time $t_{\text{wait}}$ and the detuning voltage for Q1-Q2 (d), Q2-Q3 (e) and Q3-Q4 (f). The initial states of two qubits (before the SWAP oscillations) are denoted on the top, and the qubit pair that is read out is indicated by the dashed arrow showing the readout pulse direction. g, The quantum circuit used to create a generalized Bell state between Q1 and Q2 and to characterize it via quantum state tomography (QST). h, Measured two-qubit density matrix of Q1-Q2, after removal of SPAM errors and using maximum-likelihood estimation (MLE). i, State fidelities and concurrence estimated from the density matrices of the Bell states of Q1-Q2, Q2-Q3 and Q3-Q4. The data of panels d-f and h-i is measured at $B=5$ mT.
Using the $\sqrt{X}$ and $\sqrt{Y}$ gates, we perform randomized benchmarking (RB) to obtain the average gate fidelities (see Fig. 2h). All 4 qubits yield average gate fidelities at or above 99%, extracted from the Clifford gate fidelities. We also measure the $\sqrt{X}$ and $\sqrt{Y}$ gate fidelities with gate set tomography (GST), of which the results are summarized in Fig. 2i. Overall, most of the fidelities in GST results are slightly lower than those from RB. These fidelity differences may stem from the presence of low-frequency noise in our system, which causes different uncertainties in the fidelity estimates between GST and RB  [55]. Full details on the quantum process for those two gates, derived from GST, are given in Extended Data Fig. 7 and Supplementary Information section VIII.
@@ IV Two-qubit gate
where $ij$ and $kl$ refer to the respective qubit dot pair, and the interqubit coupling $J_{\text{coup}}=(J_{ik}+J_{jl})/2$. The coupling term is reminiscent of two well-known interaction Hamiltonians. If the factor $1/2$ of the $\sigma_{z}\sigma_{z}$ coupling term were 1 instead, we recover the exchange Hamiltonian that generates the SWAP gate and the universal $\sqrt{\text{SWAP}}$ gate. If that factor were zero, only the flip-flop terms would survive, which generate the iSWAP and $\sqrt{\text{iSWAP}}$ gate. The coupling Hamiltonian in Eq. 2 thus generates a SWAP-style gate that is not a standard two-qubit gate but is also universal from the perspective of quantum computing (see Supplementary Information section VII). For simplicity, we call it a SWAP gate in the remainder of this work.
@@
Fig. 3d-f show the resulting SWAP oscillations for Q1-Q2, Q2-Q3, and Q3-Q4. Chevron-style patterns are observed with the energies of the two qubits aligned in the middle of the patterns. Moving away from the middle, the energy of one qubit is shifted with respect to that of the other. This qubit-qubit energy detuning tilts the rotation axis and accelerates the rotation. Looking closely, the chevron patterns are not symmetrical. This can be understood by the fact that the qubit energy does not vary linearly with detuning. In some panels, single-qubit oscillations around an axis close to the $x$-axis are also observed, such as the data at $\varepsilon_{26}=4$ mV in Fig. 3e. These $\varepsilon_{ij}$ values are close to zero interdot detuning, and when $J(\varepsilon_{ij}=0)$ is not much larger than $\overline{E}_{z}$, such rotations are expected. We note that SWAP oscillations between |$\ket{ST_{-}}$ and |$\ket{T_{-}S}$ were also observed in previous research on simulating the dynamics of an antiferromagnetic spin chain and resonating-valence-bond states based on the Heisenberg model in four-quantum-dot systems [18, 19].
With a combination of single-qubit gates and the SWAP-style gate, we prepare a Bell state and characterize it using quantum state tomography (QST). The pulse sequence shown in Fig. 3g is expected to produce a generalized Bell state |$\ket{\psi}=\frac{1}{\sqrt{2}}(\ket{ST_{-}}+e^{i\theta}\ket{T_{-}S})$ between the two qubits, where $\theta$ is a single-qubit phase term, followed by single-qubit gates ($I$, $\sqrt{X}$, $\sqrt{Y}$) applied to both qubits to achieve basis changes before measurement along the $z$-axis. The density matrix of the Bell state formed by Q1-Q2 is shown in Fig. 3h, and the acquired fidelity and concurrence of all the neighbouring qubit pairs are shown in Fig. 3i (see Extended Data Fig. 9 for details). The fidelities are in the range of 73(1)-90(1)and the concurrence ranges from 21(4)to 64(4)%. Concurrence is a measure of the entanglement between two qubits, which ranges from 0 (no entanglement) to 1 (maximal entanglement). Therefore, the measured concurrence for all the qubit pairs demonstrates the implemented $\sqrt{\text{SWAP}}$ gate can generate entanglement between qubits. To evaluate the performance of the $\sqrt{\text{SWAP}}$ gate, we perform two-qubit GST on Q1 and Q2. By fitting the result to a theoretical model, a gate fidelity of $\sim$80is obtained (see Extended Data Fig. 9 and Supplementary Information section VIII for details).
@@
Figure 4: Implementation of a quantum circuit for entanglement generation and quantum state transfer. a, Quantum circuit with all the qubits initialized into the singlet. An $X$ gate (32 ns) rotates Q2 into a triplet state, and a SWAP interaction for a variable time $t_{\text{SWAP}}$ periodically produces entanglement between Q1 and Q2. Two subsequent SWAP gates (30 ns and 24 ns) transfer the state of Q2 to Q4 and a final single-qubit rotation of Q4 for a variable time $t_{\text{Q4}}$ is followed by Q4 readout. The delay time between each quantum gate is set to zero. b,c, Experimental (b) and numerical (c) results after running the quantum circuit of panel a, with triplet probabilities $P_{T}$ of Q4 shown as a function of $t_{\text{SWAP}}$ and $t_{\text{Q4}}$. d, Linecuts from a showing triplet probabilities $P_{T}$ of Q4 as a function of control time $t_{\text{Q4}}$. The data is vertically shifted by 0.5 for clarity. All the data above are measured at $B=10$ mT.
@@ V Quantum circuit implementation
Finally, using a combination of the single- and two-qubit gates demonstrated above, we aim to implement a quantum circuit designed to create and distribute an entangled state across the array. As shown in Fig. 4a, we first initialize Q1 and Q2 into |$\ket{ST_{-}}$ by applying a $\pi$ rotation on Q2 and then activate a SWAP interaction between Q1 and Q2 for a duration $t_{\text{SWAP}}$. This interaction is expected to generate entanglement when $t_{\text{SWAP}}$ corresponds to a quarter period, i.e. for a $\sqrt{\text{SWAP}}$ gate. Next, we apply consecutive half-period SWAP gates of Q2-Q3 and Q3-Q4 to transfer the state of Q2 to Q4 via Q3. Finally, we perform a single-qubit $x$-axis rotation of Q4 for a time $t_{\text{Q4}}$ and measure Q4.
The experimental results are shown in Fig. 4b, where the single-qubit oscillations of Q4 as a function of $t_{\text{Q4}}$ are modulated in phase by $t_{\text{SWAP}}$, resulting in a checkerboard pattern. The underlying mechanism is that the state of Q2 oscillates as a function of $t_{\text{SWAP}}$, as quantum information is periodically exchanged between Q1 and Q2. Therefore the state of Q4 following the quantum state transfer also oscillates with $t_{\text{SWAP}}$. Where the evolution of Q4 changes phase, $t_{\text{SWAP}}$ corresponds to the duration of a $\sqrt{\text{SWAP}}$ operation (modulo an integer number of SWAP operations), at which point maximal entanglement between Q1 and Q2 is expected. When two qubits are maximally entangled, the density matrix of each qubit by itself is fully mixed. At this point, the measured $P_{T}$ of Q4 should not oscillate as a function of $t_{\text{Q4}}$. This is indeed what we observe, see Fig. 4d, where we show the linecuts from Fig. 4b. A trace without apparent oscillations is observed between two sets of out-of-phase oscillations of Q4, as expected. The same features are seen in Fig. 4c, which shows the ideally expected checkerboard pattern obtained from numerical simulations of the protocol of Fig. 4a, assuming perfect initialization, operations and readout.
@@
We also characterize the remote Bell state of Q1 and Q4 by performing QST. The experiment was performed at $B=5$ mT and the quantum circuit is similar except Q1 was initialized into a triplet instead of Q2. The resulting Bell state fidelity is 75(2)and the concurrence is 22(4)%. Compared to the concurrence of the Bell state of Q1-Q2 before state transfer, which is 90(1)%, the remote entanglement is reduced by the transfer process of two consecutive SWAP gates.
@@ VI Conclusion
In conclusion, we have experimentally demonstrated initialization, readout, and universal control of four singlet-triplet ($S-T_{-}$) qubits in a 2x4 germanium quantum dot array. By using randomized benchmarking and quantum state tomography, we obtain average single-qubit gate fidelities of 99.49(8)-99.84(1)and Bell state fidelities of 73(1)-90(1)for all the nearest qubit pairs. For the $\sqrt{\text{SWAP}}$ gate, we estimate a gate fidelity of $\sim$80by fitting the GST result to a theoretical model. Furthermore, through independent control of the exchange interactions between any pair of neighbouring spins across the device, we are able to implement a quantum circuit that spans the entire array, yielding remote entanglement of two singlet-triplet qubits with a Bell state fidelity of 75(2)and a concurrence of 22(4)%.
With four universally controlled qubits in a bilinear array, these results put baseband-controlled singlet-triplet spin qubits in germanium firmly on the map as a potential candidate for large-scale quantum computing. In future experiments, the two-qubit gate fidelity must be increased in order to allow fault-tolerant quantum computation. The gate fidelities can be potentially improved by suppressing low-frequency noise using feedback control or pulse optimization [28, 47, 56] and by a more detailed modeling of the effects of spin-orbit interaction. Additionally, other types of two-qubit gates like the CZ gate for $S-T_{-}$ qubits can be investigated. Additional improvements can be reached if the tunnel barriers are more tunable, which can be achieved by depositing the barrier gates either before [10, 19] or together with  [57] the plunger gates. Moreover, with programmable control of exchange interactions in the array, this spin ladder can also be used for analog simulation of a wealth of rich physical phenomena such as quantum magnetism [58].